(a) Consider f : R → R, f(x) = x
2
and g : R → R, g(x) = x + 5.
(i) What is g â—¦ f?
(ii) What is f â—¦ g
1
Expert's answer
2020-08-17T18:58:38-0400

"\circ" (circle sign) denotes here a function composition.

(i) "g \circ f = g(f(x)) = g(x^2) = x^2 +5"

(ii)"f \circ g = f(g(x)) = f(x+5) = (x+5)^2 = x^2 + 10x+25"